OpenStudy (anonymous):

if n=5^2000+5^2002, then what are the prime factors of n? a. 5 only b. 2 and 5 only c. 2,5, and 10 only d. 2,5, and 13 only e. 2,5,1000, and 1001 only

jimthompson5910 (jim_thompson5910):

Factor out the GCF 5^2000 to get n=5^2000+5^2002 n=5^2000+5^(2000+2) n=5^2000+5^2000*5^2 n=5^2000 * ( 1 +5^2 ) n=5^2000 * ( 1 +25 ) n=5^2000 * ( 26 ) n=5^2000 * ( 2*13 ) n=2 * 5^2000 * 13